# Generative Retrieval with Large Language Models

### v2

@@ Abstract
When completing knowledge-intensive tasks, humans sometimes need not just an answer but also a corresponding reference passage for auxiliary reading. Previous methods required obtaining pre-segmented article chunks through additional retrieval models. This paper explores leveraging the parameterized knowledge stored during the pre-training phase of large language models (LLMs) to independently recall reference passage from any starting position. We propose a two-stage framework that simulates the scenario of humans recalling easily forgotten references. Initially, the LLM is prompted to recall document title identifiers to obtain a coarse-grained document set. Then, based on the acquired coarse-grained document set, it recalls fine-grained passage. In the two-stage recall process, we use constrained decoding to ensure that content outside of the stored documents is not generated. To increase speed, we only recall a short prefix in the second stage, and then locate its position to retrieve a complete passage. Experiments on KILT knowledge-sensitive tasks have verified that LLMs can independently recall reference passage location in various task forms, and the obtained reference significantly assists downstream tasks.
Generative Retrieval with Large Language Models
Ye Wang1, Xinrun Xu2,3, Rui Xie4, Wenxin Hu1, Wei Ye4 1East China Normal University, 2University of Chinese Academy of Science 3Institute of Software Chinese Academy of Science, 4Peking University yewang@stu.ecnu.edu.cn, xuxinrun20@mails.ucas.ac.cn, ruixie@pku.edu.cn wxhu@cc.ecnu.edu.cn, wye@pku.edu.cn
@@ 1 Introduction
Figure 1: The common process of human recall of easily forgotten information often involves an intermediate step, first recalling easily memorable documents, then recalling to find the passage.
Knowledge-intensive tasks, including open-domain question answering, dialogues, and fact verification, require access to a vast amount of world or domain-specific knowledge (Petroni et al., 2021). Common methods involve utilizing external knowledge sources such as Wikipedia and employing additional sparse or dense retrieval models to first retrieve relevant passage from Wikipedia as reference, then predict answers under the condition of the reference (Karpukhin et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021). However, traditional sparse lexical retrieval lacks deep semantic understanding, and in dual-tower dense retrieval models, the representations of questions and passages are usually obtained independently (Karpukhin et al., 2020), leading to only shallow interactions being captured (Khattab et al., 2021).
Recently, generative retrieval methods that leverage the generative capabilities of models (Cao et al., 2021; Tay et al., 2022; Bevilacqua et al., 2022; Wang et al., 2022; Lee et al., 2022) have gained increasing attention. They generate passage identifiers through autoregressive generative models, performing deep token-level cross-attention, and interacting with the entire parameter space of models trained on the target corpus to overcome bottlenecks. These retrieval methods have proven effective across various domains and tasks. However, both dense retrieval and previous generative retrieval methods can only obtain predefined and segmented passages (e.g., 100 words), making it difficult to freely choose the starting point of the reference. This somewhat affects the flexibility of obtaining references and the naturalness of reading. Humans, on the other hand, can recall any seen passage from memory and naturally recall it from any starting point.
Large language models (LLMs) have proven to possess strong generative understanding capabilities that can generalize to multiple tasks through prompting, and they have parameterized memory of extensive document knowledge during the pre-training phase. We explore whether LLMs can independently recall seen passage as reference from any starting position through prompting. To this end, we mimic the human process of recalling references: first recalling related coarse-grained documents or pages, then recalling fine-grained reference passage, as shown in Figure 1. Specifically, we employ a two-stage framework. In the first stage, we prompt the large language model (LLM) to recall document title identifiers, using a prefix tree (Trie) (Cormen et al., 2022) to constrain decoding and ensure the recalled titles exist within the knowledge source. In the second stage, we construct a new FM-index (Ferragina and Manzini, 2000) with the high-ranking document set obtained from the first stage, which can effectively identify any substring within the set, ensuring that we can start constrained generation from any location in the document. We prompt the LLM to recall the required reference passage under this FM-index constraint. To leverage information from both stages of recall, we ultimately use a weighted score from the two-stage constrained decoding to select the final reference.
Although LLMs have impressive capabilities, generating complete reference passage may be too time-consuming. To address this issue, we explore whether LLMs can locate passage by recalling a prefix only. First, we prompt the LLM to generate a short prefix, then use the FM-index to locate the document containing this prefix within the document set. Finally, we use the Knuth-Morris-Pratt (KMP) algorithm to identify the starting position of the prefix in the document, determining a longer passage as the final reference from this position. We find that this method significantly speeds up the LLM’s recall of reference. And the location of such a short prefix still maintains a high quality of reference retrieval. Our framework requires no additional retrieval models and pre-segmentation, allowing the LLM to independently and naturally obtain natural reference from articles of any length, ultimately naming it LLM2GR.
We conduct comprehensive experiments on 6 knowledge-sensitive tasks from the KILT benchmark (Petroni et al., 2021). We find that in a zero-shot setting, merely prompting the open-source LLMs Llama (Touvron et al., 2023a) and Llama-2 (Touvron et al., 2023b) enables the effective recall of needed page documents and fine-grained reference passage under the LLM2GR framework, enhancing the performance of downstream tasks. In summary, our main contributions are as follows:
  • We explore for the first time the capability of LLMs to recall fine-grained reference passage under constraints with prompting.
  • We propose a two-stage framework that mimics the human information search process, first recalling coarse-grained documents, then recalling fine-grained reference.
  • By merely recalling a short prefix and locating it, we significantly speed up the recall of reference and verify the ability of LLMs to recall and locate.
  • Across 6 knowledge-sensitive tasks, our framework excels in page and passage-level evaluations and can significantly improve the performance of downstream tasks.
@@
Figure 2: Shows the LLM2GR framework. First, all Wikipedia titles are stored in a prefix tree, then the LLM is prompted to recall title identifiers under this prefix tree constraint. Subsequently, an FM-index is constructed from the top $k$ documents obtained, and the LLM recalls reference passage under the new constraint. The grey area represents the short prefix, used for the next step of locating and extracting the complete passage.
2 LLM2GR
@@
In this section, we detail our two-stage framework, LLM2GR. In the first stage, we prompt the LLM to recall title identifiers, which serve as candidate documents for the next stage. In the second stage, the LLM is prompted to recall reference passage from the documents obtained in the first stage. To increase speed, we only recall a short prefix, then locate and extract the reference within the document. The structure is shown in Figure 2.
2.1 First Stage: Coarse-Grained Document Recall
When facing knowledge-sensitive tasks, due to the difficulty of directly recalling fine-grained reference passage, as seen in the experimental analysis of directly recalling fine-grained passage in subsection 3.3, we can divide the recall process into two stages. Initially, we obtain a set of documents that are easier to recall, such as Wikipedia pages. For Wikipedia page documents with titles, we can directly prompt the LLM to recall existing title identifiers, uniquely determining the title through constrained generation. This method leverages the knowledge stored within the LLM and can also ensure that the generated title is uniquely existing.
@@ 2.1 First Stage: Coarse-Grained Document Recall
We define the set of all Wikipedia titles as $T$ and the set of all documents as $D$, such that each title and document uniquely correspond to each other. First, we store all Wikipedia titles $T$ in a prefix tree (Trie) (Cormen et al., 2022). At each step in the autoregressive generation of the LLM, we determine the set of tokens to be generated next by using both the prefix tree and the previously generated tokens, masking the logits of tokens not belonging to this set as $\infty$- ∞. In other words, the prefix tree acts as a navigation structure, guiding the model to generate tokens along a path corresponding to a known title $t$ in the set $T$. A detailed introduction to Trie can be found in Appendix A.1. We compute the score of autoregressive generation through the default implementation in the (Wolf et al., 2020) library, with the score of title $t$ given ${prompt}_{t}(x)$ as:
@@
where $y_{t}$ represents the set of tokens in the title $t$, $l_{t}$ and $|y_{t}|$ represent the number of tokens generating the title identifier, $\theta$ is the model’s parameters.
2.2 Second Stage: Fine-Grained Passage Recall
To enable LLMs to freely recall reference passage from any starting position, we utilize the FM-index constraint (Ferragina and Manzini, 2000). The FM-index can be considered a specialized prefix tree that supports searching from any position. Given a starting token or string, the FM-index can provide a possible list of token successors within $O(Vlog(V))$ time, where $V$ is the size of the vocabulary. A detailed introduction to the FM-index can be found in Appendix A.2. Here, the FM-index can be pre-built for all documents, ready to be retrieved when needed to omit the time for additional construction.
@@ 2.3 Short Prefix Recall and Localization
Although LLMs can effectively recall longer reference passage directly, their expensive inference speed somewhat diminishes their application value. In fact, generating a passage of about 150 to 200 tokens requires a significant amount of computational resources and time, which is unacceptable in many real-world scenarios. The length of generation is a key factor hindering its speed. To address this challenge, we explore whether LLMs can locate reference passage by merely recalling a prefix. For this, we propose the method Short Prefix Recall Location (SPRL). The basic idea of SPRL is to first prompt the generation of a relatively short segment prefix, then locate and extract the complete passage containing this prefix within the source document set. Specifically, we divide the entire process into the following two main steps.
Initially, given an input question $q$, we prompt the LLM to recall a short text prefix $p_{s}$, consisting of $l_{p_{s}}$ tokens, guided by the prompt $prompt_{p}$, which is the same as used in the second stage. In this step, we set $l_{p_{s}}$ significantly smaller than the full length of the long passage, thereby saving significant generation time and computational resources.
Subsequently, we find the document $d$ corresponding to $p_{s}$ within the document set $D_{k}$ obtained in the first stage. Since we have controlled the number of documents in the document set $D_{k}$ obtained in the first stage, in most cases, we can obtain a unique document $d$ containing $p_{s}$. When we cannot obtain a unique document, we simply default to selecting the first document. In the next step, we use the Knuth-Morris-Pratt (KMP) string matching algorithm to quickly determine the starting position $st$ of $p_{s}$ in $d$, then extract a complete reference $p_{final}=d[st:st+l_{p}]$ starting from $st$, with a length of $l_{p}$ tokens. The autoregressive score of the short prefix is also calculated using Equation LABEL:eq2 and the final two-stage score is obtained using Equation LABEL:eq3 to select the final reference. In Experiment 3.2.1, we find that recalling just the prefix for localization still achieves good results in obtaining reference.
@@ 3 Experiments
In this section, we conduct comprehensive experiments on coarse-grained pages, fine-grained passage-level reference evaluation, and downstream tasks to validate the effectiveness of our framework. Additionally, we perform further analyses and experiments.
@@ 3.1 Experimental Setup
Datasets We conduct extensive experiments on 6 knowledge-sensitive tasks from the KILT benchmark (Petroni et al., 2021). These tasks include open-domain question answering tasks such as NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and ELI5 (Fan et al., 2019), the fact-checking task FEVER (Thorne et al., 2018), and the open-domain dialogue system WoW (Dinan et al., 2018). Due to the lack of publicly available test sets for the KILT dataset, all experiments are conducted on its validation set. For specific details about the datasets, please refer to Appendix B. We evaluate the quality of coarse-grained pages and fine-grained reference passage, as well as the enhancement of reference for downstream tasks.
Evaluation Metrics We use R-Precision as the evaluation metric for coarse-grained page-level. For fine-grained reference evaluation, since we generate passage from any position in the document rather than retrieving reference from pre-segmented passages, to standardize evaluation criteria, for datasets such as NQ, TriviaQA, and HotpotQA, we calculate the percentage of references containing at least one gold standard answer, termed Answer in Context. In other datasets, we measure the percentage of references containing at least one gold standard entity, termed Entity in Context. For downstream tasks, we use the metrics officially implemented by KILT: Exact Match (EM) for NQ, TriviaQA, and HotpotQA; Rouge-L for ELI5; accuracy for FEVER; and F1 for WoW.
Baseline Models We compare with several traditional retrieval models. These models all use the passage segmentation from the official KILT as the source for obtaining reference. For unsupervised retrieval models, we compare the traditional sparse retrieval model BM25111We implement BM25 retrieval using the https://github.com/castorini/pyserini repository (Robertson et al., 2009), and the dense retrieval model Contriever (Izacard et al., 2022). We also compare with the dense retrieval model DPR [^2] We conduct experiments with the trained DPR model and preprocessed vector index from the https://github.com/facebookresearch/KILT repository. (Karpukhin et al., 2020) that has been fine-tuned on the full dataset. We input the first passage retrieved by the model as the reference context into the LLM, which then reads the relevant reference to answer downstream tasks.
@@
DPR 54.74 45.68 25.46 $\mathbf{16.19}$ 56.61 26.62
LLM2GR(Llama-7b) 54.46 $\mathbf{57.03}$ 44.56 15.13 76.57 52.91
LLM2GR(Llama-13b) 54.42 55.53 46.30 12.94 77.55 34.51
LLM2GR(Llama-2-7b) 56.33 56.43 46.20 14.60 77.29 49.61
LLM2GR(Llama-2-13b) $\mathbf{57.77}$ 54.41 $\mathbf{48.70}$ 15.00 $\mathbf{83.69}$ $\mathbf{57.63}$
Table 1: Coarse-grained page-level results, measured by R-Precision. $\star$ indicates full data training. Bold data in the table represents the best results, while underlined data indicates the second-best results.
@@
DPR $\mathbf{47.94}$ 66.60 20.29 14.40 41.22 45.38
LLM2GR(Llama-7b) 36.87 58.48 25.55 15.99 54.85 59.40
LLM2GR(Llama-13b) 37.72 60.96 26.34 14.80 55.20 50.79
LLM2GR(Llama-2-7b) 38.07 62.88 27.55 $\mathbf{16.85}$ 56.23 57.79
LLM2GR(Llama-2-13b) 40.82 $\mathbf{68.20}$ $\mathbf{30.04}$ 15.06 $\mathbf{58.42}$ $\mathbf{63.43}$
Table 2: Fine-grained passage-level results, measured by Answer in Context and Entity in Context of the top 1 reference. $\star$ indicates full data training. Bold data in the table represents the best results, while underlined data indicates the second-best results.
Implementation Details We choose the 7b and 13b versions of the open-source LLMs Llama (Touvron et al., 2023a) and Llama-2 (Touvron et al., 2023b) for reference recall. Downstream tasks use Llama-2-13b as the reading model. We merge the passage fragments from KILT into complete documents, serving as the data source for recall. The length of the complete documents is arbitrary. In the recall phase, we always use a beam search generation strategy. In the first stage of generation, the beam size is set to 15, and we construct an FM-index containing the top $k=2$ documents. In the second stage, the beam size is set to 10, the length of the short prefix is $l_{p_{s}}=16$, and we extract a token length of $l_{p}=150$ as the final reference. The weight setting for the two-stage weighted method is $\alpha=0.9$. All downstream tasks use greedy decoding. The prompts used in the experiments can be found in Appendix C. All experimental results are from a single run on Tesla A100 40G GPUs.
@@
BM25 25.84 71.49 $\mathbf{27.23}$ 20.48 77.54 14.02
DPR $\mathbf{33.49}$ 72.68 23.13 $\mathbf{20.75}$ 75.27 14.17
LLM2GR(Llama-7b) 29.78 70.18 24.61 20.60 78.10 14.47
LLM2GR(Llama-13b) 29.68 71.60 25.48 20.24 78.53 14.33
LLM2GR(Llama-2-7b) 29.89 70.04 25.55 20.50 78.04 14.48
LLM2GR(Llama-2-13b) 31.69 $\mathbf{72.94}$ 26.13 20.61 $\mathbf{78.79}$ $\mathbf{14.77}$
Table 3: Downstream task results. $\star$ indicates that the retrieval model underwent full data training. Bold data in the table represents the best results, while underlined data indicates the second-best results.

### v3

# MindRef: Mimicking Human Memory for Hierarchical Reference Retrieval with Fine-Grained Location Awareness

@@ Abstract
When completing knowledge-intensive tasks, humans sometimes need an answer and a corresponding reference passage for auxiliary reading. Previous methods required obtaining pre-segmented article chunks through additional retrieval models. This paper explores leveraging the parameterized knowledge stored during the pre-training phase of large language models (LLMs) to recall reference passage from any starting position independently. We propose a two-stage framework that simulates the scenario of humans recalling easily forgotten references. Initially, the LLM is prompted to recall document title identifiers to obtain a coarse-grained document set. Then, based on the acquired coarse-grained document set, it recalls fine-grained passage. In the two-stage recall process, we use constrained decoding to ensure that content outside of the stored documents is not generated. To increase speed, we only recall a short prefix in the second stage, and then locate its position to retrieve a complete passage. Experiments on KILT knowledge-sensitive tasks have verified that LLMs can independently recall reference passage locations in various task forms, and the obtained reference significantly assists downstream tasks.  [^1] Code is available at https://github.com/www-Ye/MindRef.
MindRef: Mimicking Human Memory for Hierarchical Reference Retrieval with Fine-Grained Location Awareness
[…]
11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
@@ 1 Introduction
Knowledge-intensive tasks rely heavily on large knowledge sources (Petroni et al., 2021). Traditional methods often use retrieval models to find relevant passages from resources like Wikipedia for tasks such as question answering Izacard and Grave (2021). However, limitations exist with both sparse (lack of semantic depth) and dense retrieval (limited interaction between question and passage representations) Khattab et al. (2021). Generative retrieval methods, leveraging models’ generative abilities for deeper interaction with knowledge sources, are gaining popularity Tay et al. (2022). However, Current retrieval methods require pre-segmented passages, limiting reference flexibility like human memory. We ask: "Can LLMs bypass chunking to recall references from any position?"
@@
Figure 1: Human recall of forgotten information often involves a two-step process: recalling memorable documents first, then locating the specific passage within.
Leveraging the capabilities of LLMs, we propose MindRef, a two-stage framework for flexible passage retrieval. Inspired by human recall, we first prompt the LLM to recall relevant document titles, guided by a Trie Cormen et al. (2022). Then, using an FM-index Ferragina and Manzini (2000) built from the retrieved documents, the LLM recalls specific passages with flexible starting points. A weighted score combines both stages for final reference selection. To enhance efficiency, MindRef retrieves passages by recalling only a short prefix. The LLM generates this prefix, which is located within the documents using the FM-index and KMP algorithm. Ultimately, the algorithm identifies a longer passage as the final reference. This approach allows LLMs to access and retrieve natural references from articles of any length without relying on additional retrieval models or pre-segmentation, offering both flexibility and efficiency.
Extensive experiments on 6 KILT benchmark tasks Petroni et al. (2021) demonstrate the effectiveness of MindRef, enabling open-source LLMs like LLaMA Touvron et al. (2023a) and LLaMA-2 Touvron et al. (2023b) to retrieve documents and passages, improving downstream task performance effectively. Key contributions include:
1) We propose MindRef, a cognitively-aligned retrieval framework that formalizes the ‘document-to-detail’ mechanism of human memory into neural architectures for the first time. 2) Breaking away from pre-chunking paradigms, our method achieves chunkless reference localization through joint Trie-FMIndex constrained decoding, enabling retrieval from arbitrary document positions. 3) The SPRL co-optimization strategy delivers 4× faster inference while maintaining 95%+ accuracy via short-prefix neural recall. 4) Extensive evaluations on 6 KILT tasks demonstrate state-of-the-art performance, with MindRef-boosted LLaMA-2 achieving 78.79% accuracy on FEVER.
2 MindRef Framework
@@
Figure 2: MindRef Framework. First, all Wikipedia titles are stored in a prefix tree, then the LLM is prompted to recall title identifiers under this prefix tree constraint. Subsequently, an FM-index is constructed from the top $k$ documents obtained, and the LLM recalls reference passage under the new constraint.
@@
In this section, we detail our two-stage framework, MindRef (Figure 2). In the first stage, we prompt the LLM to recall title identifiers, which serve as candidate documents for the next stage. In the second stage, the LLM is prompted to recall reference passage from the documents obtained in the first stage. To increase speed, we only recall a short prefix, then locate and extract the reference within the document. Detailed Prompt can be found in Subsection B.2.
2.1 Stage 1: Coarse-Grained Document Recall
Recalling fine-grained reference passages directly for knowledge-intensive tasks can be challenging (Subsection B.3). Therefore, we propose a two-stage process. First, we retrieve easily-recallable documents (e.g., Wikipedia pages) by leveraging their titles as unique identifiers. Using a Trie (Section A.1) data structure Cormen et al. (2022), we prompt the LLM to recall relevant titles, ensuring they correspond to existing pages.
@@ 2.1 Stage 1: Coarse-Grained Document Recall
This first stage focuses on efficiently retrieving a set of candidate documents before proceeding to fine-grained passage retrieval within these documents in the second stage. The score for generating title $t$ given prompt ${prompt}_{t}(x)$ is calculated using the standard implementation from the library:
@@
where $y_{t}$ represents the set of tokens in title $t$, $l_{t}$ and $|y_{t}|$ represent the number of tokens used to generate the title, $\theta$ represents the model’s parameters.
2.2 Stage 2: Fine-Grained Passage Recall
Following the initial identification of relevant Wikipedia pages, we move to fine-grained passage retrieval within those documents. We employ the FM-index ((Section A.2)) constraint Ferragina and Manzini (2000), a space-efficient data structure enabling fast substring search and supporting retrieval from arbitrary positions.
After obtaining the top $k$ titles and their documents ($D_{k}$), we construct a targeted FM-index specifically for $D_{k}$, reducing the search space. These FM-indexes are pre-built for all documents to avoid on-the-fly construction.
@@ 2.3 Short Prefix Recall and Localization
While LLMs excel at recalling long passages, their inference speed hinders practical application. To address this, we propose Short Prefix Recall Location (SPRL), aiming to locate passages by recalling only a short prefix.
Initially, given a question $q$, SPRL prompts the LLM to generate a short prefix $p_{s}$ of length $l_{p_{s}}$ using the same prompt ($prompt_{p}$) as in the second stage, significantly reducing generation cost. Subsequently, SPRL attempts to identify the document $d$ containing $p_{s}$ within the document set $D_{k}$ obtained in the first stage. Due to the limited size of $D_{k}$, $p_{s}$ typically maps to a unique document $d$. If $p_{s}$ is found in multiple documents within $D_{k}$, the first such document encountered is selected by default, ensuring a deterministic outcome. Next, using the KMP algorithm, the first starting position $st$ of $p_{s}$ in $d$ is determined, and a complete reference passage $p_{final}=d[st:st+l_{p}]$ is extracted. The final score is calculated using Equations LABEL:eq2 and LABEL:eq3 to select the best reference. Experiments (Section 3.2.1) demonstrate that recalling only the prefix for localization yields effective results.
@@ 3 Experiments
In this section, we conduct comprehensive experiments on coarse-grained pages, fine-grained passage-level reference evaluation, and downstream tasks to validate the effectiveness of our framework. Additionally, we perform further analyses and experiments in the Appendix through Further Analysis, and Case Studies.
@@ 3.1 Experimental Setup
Datasets: Experiments were conducted on 6 knowledge-sensitive tasks from the KILT benchmark Petroni et al. (2021) ((Section B.1)).
Evaluation Metrics: R-Precision for page-level retrieval. Answer in Context (percentage of references containing at least one gold answer) for NQ, TriviaQA, and HotpotQA. Entity in Context (percentage of references containing at least one gold entity) for other datasets. Downstream task metrics followed the official KILT implementations.
Baseline Models: We compare with several traditional retrieval models. These models all use the passage segmentation from the official KILT as the source for obtaining reference. For unsupervised retrieval models, we compare the traditional sparse retrieval model BM25222We implement BM25 retrieval using the https://github.com/castorini/pyserini repository (Robertson and Zaragoza, 2009), and the dense retrieval model Contriever (Izacard et al., 2022). We also compare with the dense retrieval model DPR [^3] We conduct experiments with the trained DPR model and preprocessed vector index from the https://github.com/facebookresearch/KILT repository. (Karpukhin et al., 2020) that has been fine-tuned on the full dataset. We input the first passage retrieved by the model as the reference context into the LLM, which then reads the relevant reference to answer downstream tasks.
Implementation Details: LLaMA Touvron et al. (2023a) and LLaMA-2 Touvron et al. (2023b) (7b and 13b) were used for reference recall. LLaMA-2-13b served as the reading model for downstream tasks. We merge the passage fragments from KILT into complete documents, serving as the data source for recall. The length of the complete documents is arbitrary. In the recall phase, we always use a beam search generation strategy. In the first stage of generation, the beam size is set to 15, and we construct an FM-index containing the top $k=2$ documents. In the second stage, the beam size is set to 10, the length of the short prefix is $l_{p_{s}}=16$, and we extract a token length of $l_{p}=150$ as the final reference. The weight setting for the two-stage weighted method is $\alpha=0.9$. All downstream tasks use greedy decoding. The prompts used in the experiments can be found in Appendix B.2. Experiments were conducted on Tesla A100 40G GPUs.
Table 1: Coarse-grained page-level results (R-Precision). $\star$ denotes full data training. Cyan indicates best results, and pink indicates second-best.
@@
DPR 54.74 45.68 25.46 16.19 56.61 26.62
MindRef (LLaMA-7b) 54.46 57.03 44.56 15.13 76.57 52.91
MindRef (LLaMA-13b) 54.42 55.53 46.30 12.94 77.55 34.51
MindRef (LLaMA-2-7b) 56.33 56.43 46.20 14.60 77.29 49.61
MindRef (LLaMA-2-13b) 57.77 54.41 48.70 15.00 83.69 57.63
Table 2: Fine-grained passage-level results (Answer/Entity in Context for top-1 reference).
@@
DPR 47.94 66.60 20.29 14.40 41.22 45.38
MindRef (LLaMA-7b) 36.87 58.48 25.55 15.99 54.85 59.40
MindRef (LLaMA-13b) 37.72 60.96 26.34 14.80 55.20 50.79
MindRef (LLaMA-2-7b) 38.07 62.88 27.55 16.85 56.23 57.79
MindRef (LLaMA-2-13b) 40.82 68.20 30.04 15.06 58.42 63.43
Table 3: Downstream task results.
@@
BM25 25.84 71.49 27.23 20.48 77.54 14.02
DPR 33.49 72.68 23.13 20.75 75.27 14.17
MindRef (LLaMA-7b) 29.78 70.18 24.61 20.60 78.10 14.47
MindRef (LLaMA-13b) 29.68 71.60 25.48 20.24 78.53 14.33
MindRef (LLaMA-2-7b) 29.89 70.04 25.55 20.50 78.04 14.48
MindRef (LLaMA-2-13b) 31.69 72.94 26.13 20.61 78.79 14.77